OpenStudy (anonymous):

5. Given the quadratic equations 2x^2+px+9=0 and 3x^2-4x+q=0 have the same roots, find the value of p and of q. @ganeshie8

OpenStudy (triciaal):

let the roots be r and s First equation factors of ac that add up to p means factors of 18 and r + s = p second equation factors of 3q that add up to -4 means r + s= -4

OpenStudy (michele_laino):

please note that: for the equation: \[2x ^{2}+px+9=0\] I can write: \[x _{1}+x _{2}=-\frac{ p }{ 2 },x _{1}*x _{2}=\frac{ 9 }{ 2 }\] where x_1 and x_2 are the roots. whereas for the second equation, \[3y ^{2}-4y+q=0\] we can write: \[y _{1}+y _{2}=-\frac{ 4 }{ 3 },y _{1}*y _{2}=\frac{ q }{ 3 }\] where y_1 and y_2 are the relative roots. Now if we want that roots has to be equal then we can require that also the above systems have to be equal so?

OpenStudy (michele_laino):

oops...\[y _{1}+y _{2}=\frac{ 4 }{ 3 }\] sorry!

OpenStudy (asnaseer):

This one is actually much simpler - just divide the first equation by 2 and the second equation by 3 and then compare coefficients

OpenStudy (perl):

@asnaseer because they have the same roots , thats why after you divide you can set the coefficients equal to each other. equivalently, a quadratic polynomial with a coefficient of 1 for the x^2 term and which has roots r1,r2 is unique.

OpenStudy (triciaal):